OpenStudy (anonymous):

(4a^2bc^2)(-ab^3c^2)(3a^b^6c) simplify

OpenStudy (anonymous):

a

OpenStudy (anonymous):

scratch that the answer is b

OpenStudy (anonymous):

can you simplify this

OpenStudy (anonymous):

\[-12 a^{3+b^6} b^4 c^5 \]

OpenStudy (anonymous):

is this your question?\[\huge(4a^2bc^2)(-ab^3c^2)(3a^{b^{6c}})\]

OpenStudy (anonymous):

or this? \[\huge(4a^2bc^2)(-ab^3c^2)(3a^{b^6}c)\]

OpenStudy (anonymous):

@alghelp can you help us about what your question is..

OpenStudy (anonymous):

or you forgat number for a \[\huge(4a^2bc^2)(-ab^3c^2)(3a^?b^6c)\]

OpenStudy (anonymous):

From Mathematica: (4 a^2 b*c^2) (-a*b^3 c^2) (3 a^b^6*c) =\[\left(4 a^2 b c^2\right)\left(-a b^3 c^2\right)\left(3 a^{b^6} c\right) \]\[\left(-4 a^3 b^4 c^4\right)\left(3 a^{b^6} c\right) \]\[-12 a^{3+b^6} b^4 c^5 \]